hi I am having trouble finding out which formula to use for this problem can some one pleas help thanks in advance.
If the lesser of two consecutive even integers is tripled, the result is 20 more than twice the grater integer. find the two integers

Answer by mananth(16946)

If the lesser of two consecutive even integers is tripled, the result is 20 more than twice the grater integer. find the two integers
let the numbers be x & x+2
3x=20+2(x+2)
3x=20+2x+4
3x-2x=24
x=24
The integers are 24 & 26